OpenStudy (anonymous):

What are the explicit equation and domain for an arithmetic sequence with a first term of 5 and a second term of 2?

OpenStudy (anonymous):

@Miracrown @joquez14 @They_Call_Me_Narii @bre4h

OpenStudy (anonymous):

Will give Medal.

OpenStudy (anonymous):

Also fan.

OpenStudy (anonymous):

@Luigi0210 @LilliVarner99 @LexiLuvv2431 @aaronq @brittanyirvinn @ChloBeiiar15 @Destinyyyy @emilyparker1 @Fredo420 @Gabby_Taylor1224 @Hero

OpenStudy (anonymous):

@iiAriana @joannacaro123 @Keigh2015

OpenStudy (anonymous):

I'm so sorry but I haven no idea. I wish I could help hun. best of luck to youthough :)

OpenStudy (keigh2015):

Same with me.

OpenStudy (anonymous):

lol well thanks anyways!(:

OpenStudy (anonymous):

@Liyah15 @ms.hi @nerlineg @OpenSessame @poopsiedoodle @Rahammii @sunrise14 @torresa @undeadknight26

OpenStudy (anonymous):

Sorry im not go at things like that

OpenStudy (anonymous):

Domain is \[\infty > 5\]

OpenStudy (anonymous):

excuse me

OpenStudy (anonymous):

\[\infty <\]

OpenStudy (anonymous):

the choices are an = 5 − 2(n − 1); all integers where n ≥ 1 an = 5 − 2(n − 1); all integers where n ≥ 0 an = 5 − 3(n − 1); all integers where n ≥ 1 an = 5 − 3(n − 1); all integers where n ≥ 0

OpenStudy (anonymous):

Plug and chug is the best way to do this

OpenStudy (anonymous):

the answer should be C